OpenStudy (haleyelizabeth2017):

A 22-newton force acting at 48 degrees and a 65-newton force acting at 24 degrees act concurrently on an object. What is the magnitude and direction of the resultant vector?

OpenStudy (anonymous):

\[ F_x = 22\cos(48^\circ) + 65\cos(24^\circ)\\ F_y = 22\sin(48^\circ) + 65\sin(24^\circ)\\ |F| = \sqrt{F_x^2+F_y^2},\quad \theta = \arctan\left(\frac{F_y}{F_x}\right) \]

OpenStudy (haleyelizabeth2017):

Would it be 85.6 N, 30 degrees?

OpenStudy (anonymous):

Not sure, since I never did the actual math myself.

OpenStudy (haleyelizabeth2017):

When I did that...that is the closest option to what I got

OpenStudy (anonymous):

Ok, sounds good

OpenStudy (haleyelizabeth2017):

Thanks